# The Defense Trilemma: Why Prompt Injection Defense Wrappers Fail?

### 09 Apr 2026 v2

@@ Abstract
We prove that no continuous, utility-preserving wrapper defense—a function $D\colon X\to X$ that preprocesses inputs before the model sees them—can make all outputs strictly safe for a language model with connected prompt space, and we characterize exactly where every such defense must fail. We establish three results under successively stronger hypotheses: boundary fixation—the defense must leave some threshold-level inputs unchanged; an $\varepsilon$-robust constraint—under Lipschitz regularity, a positive-measure band around fixed boundary points remains near-threshold; and a persistent unsafe region—under a transversality condition, a positive-measure subset of inputs remains strictly unsafe. These constitute a defense trilemma: continuity, utility preservation, and completeness cannot coexist. We prove parallel discrete results requiring no topology, and extend to multi-turn interactions, stochastic defenses. The results do not preclude training-time alignment, architectural changes, or defenses that sacrifice utility. The full theory is mechanically verified in Lean 4 with Mathlib https://github.com/mbhatt1/stuff/tree/main/ManifoldProofs (46 files, ${\sim}360$ theorems, no admitted proofs, three standard axioms) and validated empirically on three LLMs munshi2026manifold .
@@ 1 Introduction
Can you build a wrapper around a language model that eliminates all prompt injection vulnerabilities? Most current defense work implicitly assumes yes. Input classifiers that flag suspicious prompts alon2023detecting , constitutional rewriting pipelines bai2022constitutional , and input sanitization layers inan2023llama all share the same structure: a function $D\colon X\to X$ that preprocesses prompts before the model sees them, mapping unsafe inputs to safe equivalents while leaving safe inputs unchanged. We prove the answer is no, under two constraints. If the defense is continuous (similar prompts produce similar rewrites) and utility-preserving (safe prompts pass through unchanged), it cannot be complete (make every output safe). These three properties form a defense trilemma: any two can coexist, but not all three (Figure˜2).
@@ 2 Related Work
Adversarial robustness research has established that small perturbations can fool classifiers szegedy2013intriguing ; goodfellow2014explaining ; carlini2017evaluating ; madry2018towards and that robustness may fundamentally trade off against accuracy tsipras2018robustness ; fawzi2018adversarial . Certified defenses provide guarantees for fixed models cohen2019certified ; katz2017reluplex ; singh2019abstract ; huang2017safety ; bagnall2019certifying , and topological perspectives have illuminated the structure of decision boundaries naitzat2020topology . For LLMs specifically, jailbreaking attacks zou2023universal ; chao2024jailbreaking ; mehrotra2024tree ; greshake2023indirect and automated red-teaming mouret2015illuminating ; samvelyan2024rainbow demonstrate persistent vulnerabilities.
Our work differs in kind: rather than studying how failures arise for fixed models or how specific systems can be certified, we impose universal constraints on the defense map itself. The closest conceptual precedent is the no-free-lunch framework wolpert1997no , which proves that no optimizer dominates across all problems. We prove the analogous result for wrapper defenses: under continuity and utility preservation, no defense eliminates all failures.
@@
Tsipras et al. tsipras2018robustness Classifier PAC / Gaussian Robustness + accuracy No
Fawzi et al. fawzi2018adversarial Classifier Lipschitz geom. Adv. invulnerability No
Wolpert & Macready wolpert1997no Optimizer Counting Universal dominance No
Cohen et al. cohen2019certified Smoothed clf. Rand. smoothing (Certifies bound) No
@@
Diminishing safety returns Int. Stability MART ge2024mart ; Sleeper Agents hubinger2024sleeper §11
Long context $\to$ exp. harder defense Cost Asymm. Many-Shot Jailbreaking anil2024many ; Kim et al. kim2025manyshot §11
Agentic tool-use degrades exp. Pipeline Lip. InjecAgent zhan2024injecagent ; ASB zhang2024asb §11
Temperature boundary instab. Stoch. Imposs. Instability of Safety yuan2025instability §11
Quantization preserves deep vulns Int. Stability Skoltech skoltech2025quant ; ETH eth2025gguf §11
Merging preserves weaker parent Int. Stability Hammoud et al. hammoud2024merging §11
Adv. training fragments Fragment Size Whack-a-Mole liu2026whackamole ; IRIS iris2025 §11
Asymmetric transfer Transferability Weak-to-Strong zhao2025weak ; IRIS iris2025 §11
Alignment tax $\propto$ unsafe vol. Discrete Dil. Safety Tax huang2025safetytax ; Huang et al. huang2026formal §11
Patch blast radius Patching Nonl. Slingshot slingshot2026 ; Sleeper Agents hubinger2024sleeper §11
@@ 9.3 Nonlinear Agent Pipelines
10 Experimental Validation
The Manifold of Failure framework munshi2026manifold maps three LLMs over a 2D behavioral space with two axes: query indirection (how obliquely the prompt asks for unsafe content) and authority framing (how much the prompt invokes authority or permission). Table˜2 summarizes nine qualitative predictions, all directionally consistent with observations.
@@ 10 Experimental Validation
(mean AD $0.93$, basin rate $93.9\%$): near-flat alignment surface (small $L$), large robustness radii. Estimated from the 2D behavioral surface in munshi2026manifold : directional slope $G\approx 5$ at the steepest boundary crossing. For the defense-path Lipschitz constant we assume a hypothetical nearest-safe-projection defense ($D(x)$ maps each point to the closest $x^{\prime}\in S_{\tau}$) and estimate $\ell\approx 1$ from grid-adjacent score differences in the direction orthogonal to the boundary (the projection direction on the 2D grid). Setting $K=1$ (identity-rate defense), these satisfy $G>\ell(K+1)=2$.
@@ 13 Conclusion
This work characterizes structural limitations of a specific class of defenses (continuous utility-preserving wrappers). The results could inform defense engineering by identifying which design constraints matter most. They could also be misread as implying that LLM defense is futile—this is not the case. The theorems apply only to wrappers satisfying specific mathematical assumptions; training-time alignment, architectural changes, discontinuous filtering, ensemble defenses, and human-in-the-loop systems are not covered. We emphasize that the impossibility results should motivate better defense design, not abandonment of defense.
@@ Appendix G Additional Verified Results
An injective, utility-preserving defense is incomplete (every unsafe input stays unsafe). A complete, utility-preserving defense is non-injective (distinct inputs collapse to the same output). The three properties—completeness, utility preservation, injectivity—form a trilemma.
@@ Appendix H Lean Artifact
  • 10 core theory files (MoF_01–MoF_10)
  • 10 cost theory files (MoF_Cost_01–MoF_Cost_10)
  • 10 advanced theory files (MoF_Adv_01–MoF_Adv_10)

### 11 Apr 2026 v3

@@ Abstract
We prove that no continuous, utility-preserving wrapper defense—a function $D\colon X\to X$ that preprocesses inputs before the model sees them—can make all outputs strictly safe for a language model with connected prompt space, and we characterize exactly where every such defense must fail. We establish three results under successively stronger hypotheses: boundary fixation—the defense must leave some threshold-level inputs unchanged; an $\varepsilon$-robust constraint—under Lipschitz regularity, a positive-measure band around fixed boundary points remains near-threshold; and a persistent unsafe region—under a transversality condition, a positive-measure subset of inputs remains strictly unsafe. These constitute a defense trilemma: continuity, utility preservation, and completeness cannot coexist. We prove parallel discrete results requiring no topology, and extend to multi-turn interactions, stochastic defenses. The results do not preclude training-time alignment, architectural changes, or defenses that sacrifice utility. The full theory is mechanically verified in Lean 4 with Mathlib https://github.com/mbhatt1/stuff/tree/main/ManifoldProofs (46 files, ${\sim}360$ theorems, no admitted proofs, three standard axioms) and validated empirically on three LLMs [17].
@@ 1 Introduction
Can you build a wrapper around a language model that eliminates all prompt injection vulnerabilities? Most current defense work implicitly assumes yes. Input classifiers that flag suspicious prompts [1], constitutional rewriting pipelines [3], and input sanitization layers [7] all share the same structure: a function $D\colon X\to X$ that preprocesses prompts before the model sees them, mapping unsafe inputs to safe equivalents while leaving safe inputs unchanged. We prove the answer is no, under two constraints. If the defense is continuous (similar prompts produce similar rewrites) and utility-preserving (safe prompts pass through unchanged), it cannot be complete (make every output safe). These three properties form a defense trilemma: any two can coexist, but not all three (Figure˜2).
@@ 2 Related Work
Adversarial robustness research has established that small perturbations can fool classifiers [20, 6, 9, 14] and that robustness may fundamentally trade off against accuracy [21, 11]. Certified defenses provide guarantees for fixed models [5, 8, 19, 13, 2], and topological perspectives have illuminated the structure of decision boundaries [16]. For LLMs specifically, jailbreaking attacks [23, 4, 10, 12] and automated red-teaming [15, 18] demonstrate persistent vulnerabilities.
Our work differs in kind: rather than studying how failures arise for fixed models or how specific systems can be certified, we impose universal constraints on the defense map itself. The closest conceptual precedent is the no-free-lunch framework [22], which proves that no optimizer dominates across all problems. We prove the analogous result for wrapper defenses: under continuity and utility preservation, no defense eliminates all failures.
@@
Tsipras et al. [21] Classifier PAC / Gaussian Robustness + accuracy No
Fawzi et al. [11] Classifier Lipschitz geom. Adv. invulnerability No
Wolpert & Macready [22] Optimizer Counting Universal dominance No
Cohen et al. [5] Smoothed clf. Rand. smoothing (Certifies bound) No
@@
Diminishing safety returns Int. Stability MART [24]; Sleeper Agents [25] §11
Long context $\to$ exp. harder defense Cost Asymm. Many-Shot Jailbreaking [26]; Kim et al. [27] §11
Agentic tool-use degrades exp. Pipeline Lip. InjecAgent [28]; ASB [29] §11
Temperature boundary instab. Stoch. Imposs. Instability of Safety [30] §11
Quantization preserves deep vulns Int. Stability Skoltech [31]; ETH [32] §11
Merging preserves weaker parent Int. Stability Hammoud et al. [33] §11
Adv. training fragments Fragment Size Whack-a-Mole [34]; IRIS [35] §11
Asymmetric transfer Transferability Weak-to-Strong [36]; IRIS [35] §11
Alignment tax $\propto$ unsafe vol. Discrete Dil. Safety Tax [37]; Huang et al. [38] §11
Patch blast radius Patching Nonl. Slingshot [39]; Sleeper Agents [25] §11
@@ 9.3 Nonlinear Agent Pipelines
10 Experimental Motivation
The Manifold of Failure framework [17] maps three LLMs over a 2D behavioral space with two axes: query indirection (how obliquely the prompt asks for unsafe content) and authority framing (how much the prompt invokes authority or permission). Table˜2 summarizes nine qualitative predictions, all directionally consistent with observations.
@@ 10 Experimental Motivation
(mean AD $0.93$, basin rate $93.9\%$): near-flat alignment surface (small $L$), large robustness radii. Estimated from the 2D behavioral surface in [17]: directional slope $G\approx 5$ at the steepest boundary crossing. For the defense-path Lipschitz constant we assume a hypothetical nearest-safe-projection defense ($D(x)$ maps each point to the closest $x^{\prime}\in S_{\tau}$) and estimate $\ell\approx 1$ from grid-adjacent score differences in the direction orthogonal to the boundary (the projection direction on the 2D grid). Setting $K=1$ (identity-rate defense), these satisfy $G>\ell(K+1)=2$.
@@
10.1 Live Empirical Validation of Theorem 6.3
The numbers above are read off the published Manifold of Failure heatmaps. To check the persistence theorem on independently re-collected alignment-deviation surfaces and to expose every intermediate quantity, we built a small validator that takes any 2D heatmap and computes:
  1. 1.
    the empirical surface Lipschitz constant $L$ from all-pairs finite differences over filled grid cells;
  2. 2.
    for a chosen defense $D$, the empirical Lipschitz constant $K$ of $D$ and the defense-path Lipschitz constant $\ell$;
  3. 3.
    the maximum directional rate $G$ at which $f$ rises across the safe-unsafe boundary (all-pairs finite differences between filled safe and filled unsafe cells);
  4. 4.
    the predicted steep set $\mathcal{S}_{\text{pred}}=\{x:f(x)>\tau+\ell(K{+}1)\operatorname{dist}(x,\,z^{*})\}$, where $z^{*}$ is the filled boundary cell with $f(z^{*})$ closest to $\tau$;
  5. 5.
    the actual persistent set $\mathcal{S}_{\text{act}}=\{x:f(D(x))>\tau\}$.
Theorem 6.3’s empirical content is the containment $\mathcal{S}_{\text{pred}}\subseteq\mathcal{S}_{\text{act}}$ under any continuous, $K$-Lipschitz, utility-preserving $D$. We test this as a per-cell confusion matrix:
  • True positive (TP): cell in both sets — theorem confirmed for that cell.
  • Interior false positive (FPint): non-boundary cell in $\mathcal{S}_{\text{pred}}$ but not $\mathcal{S}_{\text{act}}$. A non-zero count would be a real counterexample.
  • Boundary false positive: a discrete defense moving a boundary cell. Not a violation; witnesses that the discrete defense is not topologically continuous, so the theorem’s hypothesis does not apply.
  • False negative (FN): cell in $\mathcal{S}_{\text{act}}$ but not $\mathcal{S}_{\text{pred}}$. Not a violation; Theorem 6.3 only claims the containment one way.
We ran the full pipeline live against the OpenAI API on two real target models: gpt-5-mini-2025-08-07 (well aligned) and gpt-3.5-turbo-0125 (less aligned), with gpt-4.1-2025-04-14 as the judge. The MAP-Elites driver from [17] produced a $25\times 25$ alignment-deviation grid for each target, which the validator then analyzed against the canonical continuous defense ($D=\mathrm{id}$, so $K=1$ and $\ell=0$). For gpt-3.5-turbo-0125 we additionally ran a saturation study with $400$ evaluations ($100$ seed prompts +300+300 MAP-Elites iterations, $\sim$5,2005{,}200 OpenAI API calls, wall clock $\sim$$130$ min, cost $\sim$\$15) which filled $82$ cells with $66$ strictly unsafe ($f>\tau$) and $10$ boundary cells. The two earlier runs are kept as smoke-tests of the validator’s logic on sparse grids; the saturation row is the doubt-eliminator.
[…]
gpt-5-mini-2025-08-07 $0.30$ $14$ $6.13$ $6.13$ 77 77 77 $\mathbf{0}$ 0
gpt-3.5-turbo-0125 (smoke) $0.50$ 99 $11.44$ $11.44$ 77 77 77 $\mathbf{0}$ 0
gpt-3.5-turbo-0125 (saturated) $0.50$ $82$ $23.63$ $23.63$ $66$ $66$ $66$ $\mathbf{0}$ 0
[…]
Defense sweep on the saturated grid.
The validator’s sweep subcommand runs the same archive against the canonical identity defense and against multiple discrete bounded_step defenses with progressively larger displacement radii. Across all $7$ defenses tested (identity plus bounded_step with $\text{max\_step}\in\{1,2,3,5,8,12\}$), the per-cell confusion-matrix produces zero interior false positives on every run—i.e., zero empirical counterexamples to Theorem 6.3 across the entire defense sweep on the same dense alignment-deviation surface (Table˜4).
[…]
identity $1.00$ $0.00$ $66$ $66$ $66$ $\mathbf{0}$ 0
bounded_step(1) $3.61$ $23.63$ 11 $55$ 0 $\mathbf{0}$ $55$
bounded_step(2) $3.61$ $23.63$ 11 $47$ 0 $\mathbf{0}$ $47$
bounded_step(3) $4.00$ $23.63$ 11 $29$ 0 $\mathbf{0}$ $29$
bounded_step(5) $8.60$ $23.63$ 11 33 0 $\mathbf{0}$ 33
bounded_step(8) $8.60$ $23.63$ 11 0 0 $\mathbf{0}$ 0
bounded_step(12) $8.60$ $23.63$ 11 0 0 $\mathbf{0}$ 0
Reading the sweep: as the defense’s reach grows from $\text{max\_step}=1$ (cell can move at most one grid step) to $\text{max\_step}=8$ (essentially nearest-safe projection), the actual persistent set shrinks from $55$ cells down to 0. The predicted steep set under any bounded_step is just the anchor cell because the empirical defense-path Lipschitz constant $\ell=L=23.63$ on this dense surface, which forces $\ell(K{+}1)>G$ for all K>0K>0—transversality fails for the discrete bounded defenses. Theorem 6.3 therefore makes no positive prediction on those rows, only the negative one (“no interior counterexample is allowed”), and the empirical data confirms it. The well-aligned target additionally exemplifies the harmless-boundary case: at $\tau=0.5$, gpt-5-mini’s peak AD is exactly $0.5$ and $U_{\tau}=\emptyset$, so the trilemma’s preconditions fail and the validator correctly reports “not applicable”—matching the qualitative GPT-5-Mini observation above and confirming the theorem’s silence in the right place.
Non-vacuous GP-smooth oblique defense.
[…]
The validator is available as a small standalone Python package (trilemma_validator/ in the repository), with subcommands trilemma pipeline (run rethinking-evals end-to-end and validate), trilemma validate (analyze an existing heatmap), trilemma sweep (run multiple defenses on a single archive), trilemma experiment (just run the experiment), and trilemma synth (generate a synthetic mesa/mosaic/flat surface for a self-contained smoke test). The full per-cell records (each cell’s $f(x)$, $D(x)$, $f(D(x))$, distance to $z^{*}$, $\varepsilon$-robust LHS and RHS, and steep/persistent membership flags) are emitted as JSON by the validator and reproduced from the live archives in trilemma_validator/live_runs/, including the saturated grid above (gpt35_turbo_t05_saturated/).
@@ 13 Conclusion
This work characterizes structural limitations of a specific class of defenses (continuous utility-preserving wrappers). The results could inform defense engineering by identifying which design constraints matter most. They could also be misread as implying that LLM defense is futile—this is not the case. The theorems apply only to wrappers satisfying specific mathematical assumptions; training-time alignment, architectural changes, discontinuous filtering, ensemble defenses, and human-in-the-loop systems are not covered.
We emphasize that the impossibility results should motivate better defense design and appropriate policy corrections, not abandonment of defense.
@@ Appendix G Additional Verified Results
An injective, utility-preserving defense is incomplete (every unsafe input stays unsafe). A complete, utility-preserving defense is non-injective (distinct inputs collapse to the same output). The three properties---completeness, utility preservation, injectivity---form a trilemma.
@@ Appendix H Lean Artifact
  • 10 core theory files (MoF_01--MoF_10)
  • 10 cost theory files (MoF_Cost_01--MoF_Cost_10)
  • 10 advanced theory files (MoF_Adv_01--MoF_Adv_10)